OpenStudy (anonymous):

Solve. x + 4 – 7x = 22

OpenStudy (aakashsudhakar):

x + 4 - 7x = 22 -6x = 18 -x = 3 x = -3 Check your work! x = -3 x + 4 - 7x = 22 (-3) + 4 - 7(-3) = 22 1 - (-21) = 22 22 = 22